OpenStudy (anonymous):

PLEASE HELP ME Question 1 (Multiple Choice Worth 4 points) Consider the following system of equations: y = 5x + 6 y = −x − 7 Which description best describes the solution to the system of equations? Line y = 5x + 6 intersects line y = −x − 7. Lines y = 5x + 6 and y = −x − 7 intersect the x-axis. Lines y = 5x + 6 and y = −x − 7 intersect the y-axis. Line y = 5x + 6 intersects the origin

OpenStudy (perl):

Please show me the work you have so far.

OpenStudy (anonymous):

ok I have \[5x-5+6-5=x=1\]

OpenStudy (welshfella):

how do you solve systems of equations of this form graphically?

OpenStudy (anonymous):

I think you Just put them on a graph correct

OpenStudy (welshfella):

yes - and what does a point of intersection give you?

OpenStudy (anonymous):

the coordinates

OpenStudy (welshfella):

the coordinates of the solution to the system other probabilities are the 2 equations give the same line or they are parallel - infinite solutions and no solution respectively

OpenStudy (welshfella):

* possibilities

OpenStudy (welshfella):

if you try to solve this system and get one value for x it will be the first of the 3 types

OpenStudy (welshfella):

so if you find that the answer is the first option

OpenStudy (welshfella):

As y is equal to both 5x + 6 and -x - 7 we can say 5x + 6 = -x - 7 try solving for x and see what you get

OpenStudy (welshfella):

do you need help solving this?

OpenStudy (anonymous):

I tried it and got A

OpenStudy (welshfella):

yes a is correct option